# Point modules over the universal enveloping algebras of color Lie algebras and overlapping normal elements

### 01 Apr 2026 v1

@@ Abstract.
Let $k$ be an algebraically closed field with characteristic 0. In this paper, we define the notion of an overlapping normal element of a $\mathbb{Z}$-graded $k$-algebra. This overlapping normal element gives us the two criteria for the nonexistence of a module. The first one involves a module with low Gelfand–Kirillov dimension. Also, the second one involves a truncated point module. By using the second one, we determine the set of point modules over an Artin–Schelter regular algebra obtained as the universal enveloping algebra of a color Lie algebra.
@@ 1. Introduction
In this paper, we define the notion of an overlapping normal element of a $\mathbb{Z}$-graded $k$-algebra.
@@
Let $A$ be a $\mathbb{Z}$-graded $k$-algebra and $g$ a regular homogeneous normal element of $A$ with degree $n\geq 1$. A regular homogeneous normal element $g$ is called an overlapping normal element of $A$, if there exist $u\in k^{\times}$, $x\in A_{1}$, and $y\in A_{n-1}$ satisfies the following equations;
@@
This overlapping normal element gives us some criteria for the nonexistence of a module having a relationship to a point module. Theorem 3.4 describes a condition of the nonexistence of a module with respect to the $k$-vector space dimension. Corollary 3.6 gives a criterion for the nonexistence of a module with respect to the Gelfand–Kirillov dimension. Theorem 3.4 can treat a fat point module, but it cannot treat a truncated point module. In Theorem 4.3, we obtain another criterion for the nonexistence of a truncated point module. Both Theorems 3.4 and 4.3 can treat a point module, thereby determining the structure of the set of closed points of a point scheme: For a connected $\mathbb{N}$-graded finitely presented $k$-algebra AA generated in degree $1$ and an overlapping normal element $g$ of $A$, the set of closed points of the point scheme of $A$ completely overlaps one of $A/\langle g\rangle$. In Theorem 4.5, using this result, we determine the structure of the set of closed points of the point scheme of the universal enveloping algebra of color Lie algebras.
@@
This paper is organized as follows: In Section 2, we recall some definitions and properties of the quasi-Veronese algebra defined by Mori [15], the twisted algebra defined by Zhang [28], the dehomogenization, point modules, and color Lie algebras. In Section 3, we deal with a general case. First, we define an overlapping normal element in Definition 3.1. Also, we prove Theorem 3.4, which is a criterion for the nonexistence of a $g$-torsionfree module with an overlapping normal element $g$. In Section 4, for a connected $\mathbb{N}$-graded finitely presented $k$-algebra AA generated in degree $1$ and an overlapping normal element $g$ of $A$, we prove Theorem 4.3, which is a truncated version theorem of the nonexistence of a $g$-torsionfree module with some conditions. By Theorem 4.3, we determine the set of closed points of the point schemes of the universal enveloping algebra of color Lie algebras in Theorem 4.5. At the end of this section, we show that the inverse system of the truncated point schemes of it is constant in Corollary 4.6.
@@ 2.1. Graded algebras and filtered algebras
Definition 2.4 ([28, Proposition and Definition 4.2]).
@@
Lemma 2.7 (see [12, Chapter I 4.3 Lemma 2]).
@@ 2.2. Point schemes and point modules
    A graded module $P=\bigoplus_{i\in\mathbb{Z}}P_{i}$ is called a point module over $A$ if $P$ is cyclic (i.e., $P=AP_{0}$) and satisfies the equation; $\operatorname{dim}_{k}P_{i}=\begin{cases}1&\text{for $i\geq 0$},\\ 0&\text{for $i<0$}.\end{cases}$
@@
    For each integer $d\geq 1$, a graded module $P=\bigoplus_{i\in\mathbb{Z}}P_{i}$ is called a truncated point module of length $d+1$ over $A$ if $P$ is cyclic (i.e., $P=AP_{0}$) and satisfies the equation; $\operatorname{dim}_{k}P_{i}=\begin{cases}1&\text{for $0\leq i\leq d$},\\ 0&\text{otherwise}.\end{cases}$
@@ 2.3. Color Lie algebras
    The universal enveloping algebra $U(L)$ of a finite dimensional color Lie algebra $L$ with $L_{-}=0$ has a finite Gelfand–Kirillov dimension by [14, 8.1.14] and [21, 4.C] (also, see [19]).
@@
    In particular, when a $\mathbb{Z}$-graded $k$-algebra $U(L)$ is a connected $\mathbb{N}$-graded $k$-algebra generated in degree $1$, $U(L)$ is an AS-regular algebra by [11, 6.3 Theorem] and [17, Theorems 3.1, 3.2].
In the end of this subsection, for the universal enveloping algebra $U(L)$ of a $(G,\varepsilon)$-color Lie algebra $L$, we recall a $G$-graded free resolution of the trivial module $k$ of $U(L)$ called the color Koszul resolution of $U(L)$ ([6, Theorem 3]).
@@
3. Overlapping normal elements
In this section, for a $\mathbb{Z}$-graded $k$-algebra AA, we define an overlapping normal element of $A$ and prove some properties about this.
@@ 3. Overlapping normal elements
Let $A$ be a $\mathbb{Z}$-graded $k$-algebra and $g$ a regular homogeneous normal element of $A$ with degree $n\geq 1$. A regular homogeneous normal element $g$ is called an overlapping normal element of $A$, if there exist $u\in k^{\times}$, $x\in A_{1}$, and $y\in A_{n-1}$ satisfies the following equations;
@@
If either $x$ or $y$ is a central element of $A$, then $g=0$ holds. In particular, when $A$ is commutative, $A$ has no overlapping normal element.
For $\mathbb{Z}$-graded $k$-algebra AA, in some situations, the regularity of an overlapping normal element of $A$ is not necessarily, but $u\in k^{\times}$ is sufficient. For convenience, we assume the regularity of an overlapping normal element. We show some examples of an overlapping normal element of some algebras.
@@
    Suppose that $A$ is a down-up algebra $A(2u,-u^{2},0)$ where $u\in k^{\times}$, that is,
    $A(2u,-u^{2},0):=k\langle x,y\rangle/\langle x^{2}y-2uxyx+u^{2}yx^{2},xy^{2}-2uyxy+u^{2}y^{2}x\rangle.$
    Then g=xyuyxg=xy-uyx is an overlapping normal element of $A$.
@@
    Let $A$ be a $\mathbb{N}$-graded $k$-algebra, $f$ a homogeneous element with degree $n$, and $\delta$ a graded derivation of $A$. Suppose that $B=A[X,\delta]$ is a graded Ore extension of $A$ with a degree $1$ element $X$. If $Y=\delta(f)$ is regular in $B[Y]/\langle Y-\delta(f)\rangle$ where $Y$ is a degree $n+1$ element, then $\delta(f)$ is an overlapping normal element of it. Indeed, the definition of the Ore extension gives the following equations;
@@
Theorem 3.4 is our main result in this paper. For $\mathbb{Z}$-graded $k$-algebra AA, an overlapping normal element $g$ of $A$ gives a criterion for nonexistence of a $g$-torsionfree module over $A$.
Theorem 3.4.
Let $A$ be a $\mathbb{Z}$-graded $k$-algebra, g=xyuyxg=xy-uyx an overlapping normal element of $A$ with degree $n$, and $M=\bigoplus_{i\in\mathbb{Z}}M_{i}$ a $g$-torsionfree left bounded (i.e., $M_{i}=0$ for $i\ll 0$) graded module of $A$. If there exists $c\in\mathbb{R}$ such that $\operatorname{dim}_{k}M_{s}<c$ for $s\gg 0$, then $M=0$.
@@
Let $A^{[n]}$ be the $n$-th quasi-Veronese algebra of $A$, $\mathbf{1}$ the identity of $A^{[n]}$, and $\nu$ the graded automorphism determined by $\nu(a)g=ga$ for $a\in A$. Also, the map defined by $\nu^{[n]}((a_{i,j}))=(\nu(a_{i,j}))$ for $(a_{i,j})\in(A^{[n]})_{r}$ is a graded automorphism of $A^{[n]}$ in [15]. We denote by $\mathbf{g}$ the element defined as
where $\delta_{i,j}$ is the Kronecker delta. For $\mathbf{a}=(a_{i,j})\in(A^{[n]})_{r}$, the following equation holds;
$\mathbf{ga}=\left(\sum_{l=0}^{n-1}\delta_{l,j}ga_{i,l}\right)=(ga_{i,j})=(\nu(a_{i,j})g)=\left(\sum_{l=0}^{n-1}\nu(a_{l,j})\delta_{i,l}g\right)=\nu^{[n]}(\mathbf{a})\mathbf{g}.$
Then $\mathbf{g}$ is a regular normal element in $A^{[n]}$ with degree $1$. By [3] and [28, Example 2.12], the set $\nu^{[n]}=\{(\nu^{[n]})^{i}\mid i\in\mathbb{Z}\}$ is a twisting system of $A^{[n]}$ and $\mathbf{g}$ is a central element of the twisted algebra $^{\nu^{[n]}}A^{[n]}$. Let $\mathcal{R}$ be the dehomogenization of $^{\nu^{[n]}}A^{[n]}$ with respect to $\mathbf{g}$ and $F\mathcal{R}=\{F_{i}\mathcal{R}\}_{i\in\mathbb{Z}}$ the filtration of $\mathcal{R}$ defined by the dehomogenization. We consider the first Weyl algebra $W_{1}(k)=k\langle X,Y\rangle/\langle XY-YX-1\rangle$ with the Bernstein filtration $\mathcal{B}=\left\{\operatorname*{\mathchoice{\Xdsum}{\Xdsum}{\Xtsum}{\Xtsum}}_{|\alpha|+|\beta|\leq i}c_{\alpha,\beta}Y^{\alpha}X^{\beta}\in W_{1}(k)\middle|c_{\alpha,\beta}\in k\right\}_{i\in\mathbb{Z}}$. Note that $\mathcal{B}_{i}=0$ for $i<0$ (see [8]). Let $\varphi:W_{1}(k)\rightarrow\mathcal{R}$ be the $k$-algebra homomorphism defined by mapping of generators as follows;
@@
Since $\nu(x)=u^{-1}x$, $\nu(y)=uy$, and $\nu(g)=g$, the following equations are obtained:
@@
Clearly, the homomorphism $\varphi$ preserves the filtrations $\mathcal{B}$ and $F\mathcal{R}$. Also, Lemma 2.2 and Proposition 2.5 provide the following categorical equivalence;
@@
$\operatorname{dim}_{k}F_{s}\mathcal{N}=\operatorname{dim}_{k}{}^{\nu^{[n]}}(Q(M))_{s}=\operatorname*{\mathchoice{\Xdsum}{\Xdsum}{\Xtsum}{\Xtsum}}_{i=0}^{n-1}\operatorname{dim}_{k}M_{sn-i}<cn\quad\text{ for $s\gg 0$.}$
We assume that $\mathcal{N}\neq 0$. Then there exists a good filtration $\Gamma=\{\Gamma_{i}\}_{i\in\mathbb{Z}}$ of $\mathcal{N}$ and exist $c_{1}\neq 0,c_{0}\in\mathbb{Q}$ such that $\operatorname{dim}_{k}\Gamma_{s}=c_{1}s+c_{0}$ for $s\gg 0$. Furthermore, there exists $s_{0}\in\mathbb{Z}$ such that $\Gamma_{i}\subset F_{i+s_{0}}\mathcal{N}$ for $i\geq 0$ by [8, 8.3.2]. Hence, the inequality $cn>\operatorname{dim}_{k}F_{s+s_{0}}\mathcal{N}\geq\operatorname{dim}_{k}\Gamma_{s}=c_{1}s+c_{0}$ holds for $s\gg 0$. This is a contradiction. Therefore, $\mathcal{N}=0$ and then $M=0$. ∎
@@
Let $M=\bigoplus_{i\in\mathbb{Z}}M_{i}$ be a $g$-tosionfree left bounded graded module such that $\operatorname{dim}_{k}M_{s}=c_{1}s+c_{0}$ for $s\gg 0$ where $c_{1}\neq 0$, $c_{0}\in\mathbb{Q}$ (e.g., $M$ is a line module). By the similar proof of Theorem 3.4, the dehomogenization of $^{\nu}Q(M)$ is a holonomic module of $W_{1}(k)$.
To obtain a more useful criterion as a corollary of Theorem 3.4, we recall the definition of the Gelfand–Kirillov dimension. Let $A$ be a finitely generated $k$-algebra generated by a finite dimensional $k$-vector space $1\in V$ and $M$ a finitely generated $A$-module generated by $M_{0}$. The Gelfand–Kirillov dimension of $A$ (resp. $M$) is defined as follows (see [14, 8.1.9, 8.1.11]);
$\mathrm{GKdim}(A):=\overline{\lim_{n\rightarrow\infty}}\log_{n}\operatorname{dim}_{k}V^{n}\quad(\,\text{resp. }\mathrm{GKdim}(M):=\overline{\lim_{n\rightarrow\infty}}\log_{n}\operatorname{dim}_{k}V^{n}M_{0}\,).$
Corollary 3.6.
Let $A$ be a $\mathbb{Z}$-graded finitely generated $k$-algebra generated in degree $1$, $g$ an overlapping normal element of $A$ with degree $n$, and $M$ a finitely generated $g$-torsionfree graded $A$-module. If $\mathrm{GKdim}(M)<2$, then $M=0$.
Proof.
Since $M$ is a $g$-torsionfree module, $\operatorname{dim}_{k}(M/gM)<\infty$ holds by [14, 8.3.5] and [14, 8.1.17]. For $s\geq 0$, let $p,r$ be integers satisfying $s=pn+r$ and $0\leq r<n$. Then the following inequality hold;
$\operatorname{dim}_{k}M_{s}$ $=\operatorname*{\mathchoice{\Xdsum}{\Xdsum}{\Xtsum}{\Xtsum}}\limits_{i=1}^{p}(\operatorname{dim}_{k}M_{in+r}-\operatorname{dim}_{k}M_{(i-1)n+r})+\operatorname{dim}_{k}M_{r}$
$\leq\operatorname{dim}_{k}(M/gM)+\operatorname*{\mathchoice{\Xdsum}{\Xdsum}{\Xtsum}{\Xtsum}}\limits_{j=0}^{n-1}\operatorname{dim}_{k}M_{j}.$
Therefore, by Theorem 3.4, we have the conclusion. ∎
4. Overlapping normal elements and point modules
In this section, for a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$, we show more details of a relationship between a point scheme of $A$ and an overlapping normal element of $A$. In the end of this section, we show the structure of the set of closed points of the point scheme of the universal enveloping algebra of a color Lie algebra.
4.1. Overlapping normal elements and truncated point modules
Throughout this subsection, let $A$ be a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$ unless otherwise specified. In this subsection, we show a relationship between an overlapping normal element $g$ of $A$ and a truncated $g$-torsionfree point module over $A$.
First, we recall useful equations obtained by the definition of an overlapping normal element of $A$. For an overlapping normal element g=xyuyxg=xy-uyx of $A$, the equations $x^{2}y-2uxyx+u^{2}yx^{2}=0$ and $xy^{2}-2uyxy+u^{2}y^{2}x=0$ holds. Then the same equations in a down-up algebra holds as in [10].

### 11 Apr 2026 v2

@@ Abstract.
Let $k$ be an algebraically closed field with characteristic 0. In this paper, we define the notion of a $q^{\prime}$-Heisenberg normal element of a $\mathbb{Z}$-graded $k$-algebra. This $q^{\prime}$-Heisenberg normal element gives the structure of some sets of modules related to point modules. Also, we determine the set of point modules over an Artin–Schelter regular algebra obtained as the universal enveloping algebra of a color Lie algebra. Moreover, we give a concrete integer such that the inverse system of truncated point schemes of it is constant.
@@ 1. Introduction
A module called a point module over a connected $\mathbb{N}$-graded finitely presented $k$-algebra AA generated in degree $1$ is important from the following fact ([2]): The set of closed points of the point scheme of $A$ corresponds one-to-one to the set of isomorphism classes of point modules over $A$. In this paper, in order to study a point module, we define the notion of a $q^{\prime}$-Heisenberg normal element of a $\mathbb{Z}$-graded $k$-algebra.
@@
Let $A$ be a $\mathbb{Z}$-graded $k$-algebra and $g$ a regular homogeneous normal element of $A$ with degree $n\geq 1$. A regular homogeneous normal element $g$ is called a $q^{\prime}$-Heisenberg normal element of $A$, if there exist $u\in k^{\times}$, $x\in A_{1}$, and $y\in A_{n-1}$ satisfies the following equations;
@@
The relations in the above definition of a $q^{\prime}$-Heisenberg normal element are similar to defining relations of another $q$-analog of the universal enveloping algebra of the Heisenberg algebra (cf. Kirkman–Small [13]). To emphasize a correspondence to another case, we denote $q^{\prime}$ instead of simply denoting $q$. This $q^{\prime}$-Heisenberg normal element gives us the structure of some sets of modules having a relationship to a point module. Assume that $A$ is a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$ and $g$ is a $q^{\prime}$-Heisenberg normal element of $A$. Proposition 3.7 shows that the set of base-point modules over $A$ completely overlaps one of $A/\langle g\rangle$. Also, in Theorem 4.3, we obtain a criterion for the nonexistence of a truncated point module. Therefore, we obtain the fact that some set of truncated point modules over $A$ completely overlaps one of $A/\langle g\rangle$. In Theorem 4.5, using this result, we determine the structure of the set of closed points of the point scheme of the universal enveloping algebra of color Lie algebras.
@@
This paper is organized as follows: In Section 2, we recall some definitions and properties of the quasi-Veronese algebra defined by Mori [18], the twisted algebra defined by Zhang [31], the dehomogenization, point modules, and color Lie algebras. In Section 3, we define the notion of a $q^{\prime}$-Heisenberg normal element and prove basic properties. In Proposition 3.7, we prove that the non-existence of a $g$-torsionfree base-point module for a $q^{\prime}$-Heisenberg normal element $g$. In Section 4, we prove Theorem 4.3, which is a theorem of the nonexistence of a $g$-torsionfree module with some conditions. By Theorem 4.3, we determine the set of closed points of the point schemes of the universal enveloping algebra of color Lie algebras in Theorem 4.5. At the end of this section, we show that the inverse system of the truncated point schemes of it is constant in Corollary 4.6.
@@ 2.1. Graded algebras and filtered algebras
Definition 2.4 ([31, Proposition and Definition 4.2]).
@@
Lemma 2.7 (see [15, Chapter I 4.3 Lemma 2]).
@@ 2.2. Point schemes and point modules
    A graded module $P=\bigoplus_{i\in\mathbb{Z}}P_{i}$ is called a point module over $A$ if $P=AP_{0}$ and $P$ satisfies the equation; $\operatorname{dim}_{k}P_{i}=\begin{cases}1&\text{for $i\geq 0$},\\ 0&\text{for $i<0$}.\end{cases}$
@@
    For each integer $d\geq 1$, a graded module $P=\bigoplus_{i\in\mathbb{Z}}P_{i}$ is called a truncated point module of length $d+1$ over $A$ if $P=AP_{0}$ and $P$ satisfies the equation; $\operatorname{dim}_{k}P_{i}=\begin{cases}1&\text{for $0\leq i\leq d$},\\ 0&\text{otherwise}.\end{cases}$
Moreover, we recall the definition of a base-point module.
Definition 2.10 ([7]).
Let $A$ be a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$ and $c\geq 1$ an integer. A graded module $M=\bigoplus_{i\in\mathbb{Z}}M_{i}$ is called a base-point module over $A$ if $AM_{0}, MM$ satisfies the equation;
$\operatorname{dim}_{k}M_{i}=\begin{cases}c&\text{for $i\geq 0$};\\ 0&\text{for $i<0$};\end{cases}$
and $M$ is $1$-critical with respect to the Gelfand–Kirillov dimension.
Remark that, for a base-point module $M$, when $c=1$, $M$ is a point module. Also, when $c\geq 2$, $M$ is called a fat point module ([5]).
@@ 2.3. Color Lie algebras
    The universal enveloping algebra $U(L)$ of a finite dimensional color Lie algebra $L$ with $L_{-}=0$ has a finite Gelfand–Kirillov dimension by [17, 8.1.14] and [24, 4.C] (also, see [22]).
@@
    In particular, when a $\mathbb{Z}$-graded $k$-algebra $U(L)$ is a connected $\mathbb{N}$-graded $k$-algebra generated in degree $1$, $U(L)$ is an AS-regular algebra by [14, 6.3 Theorem] and [20, Theorems 3.1, 3.2].
In the end of this subsection, for the universal enveloping algebra $U(L)$ of a $(G,\varepsilon)$-color Lie algebra $L$, we recall a $G$-graded free resolution of the trivial module $k$ of $U(L)$ called the color Koszul resolution of $U(L)$ ([8, Theorem 3]).
@@
3. $q^{\prime}$-Heisenberg normal elements
In this section, for a $\mathbb{Z}$-graded $k$-algebra AA, we define a $q^{\prime}$-Heisenberg normal element of $A$ and prove some basic properties about this.
@@ 3. $q^{\prime}$-Heisenberg normal elements
Let $A$ be a $\mathbb{Z}$-graded $k$-algebra and $g$ a regular homogeneous normal element of $A$ with degree $n\geq 1$. A regular homogeneous normal element $g$ is called a $q^{\prime}$-Heisenberg normal element of $A$, if there exist $u\in k^{\times}$, $x\in A_{1}$, and $y\in A_{n-1}$ satisfies the following equations;
@@
If either $x$ or $y$ is a central element of $A$, then $g=0$ holds. In particular, when $A$ is commutative, $A$ has no $q^{\prime}$-Heisenberg normal element.
For $\mathbb{Z}$-graded $k$-algebra AA, in some situations, the regularity of a $q^{\prime}$-Heisenberg normal element of $A$ is not necessarily, but $u\in k^{\times}$ is sufficient. For convenience, we assume the regularity of a $q^{\prime}$-Heisenberg normal element. We show some examples of a $q^{\prime}$-Heisenberg normal element of some algebras.
@@
  1. Suppose that $A(\alpha,\beta)$ is a graded down-up algebra
    $A(\alpha,\beta):=k\langle x,y\rangle/\langle x^{2}y-\alpha xyx-\beta yx^{2},xy^{2}-\alpha yxy-\beta y^{2}x\rangle.$
    where $\alpha,\beta\in k$ (see [12]). When $\alpha,\beta\neq 0$ and $\alpha^{2}+4\beta=0$ hold, $g=xy-\dfrac{\alpha}{2}yx$ is a $q^{\prime}$-Heisenberg normal element of $A$. This algebra is isomorphic to another $q$-analog of the universal enveloping algebra of the Heisenberg algebra $H^{\prime}q$ with $q=\dfrac{\alpha}{2}$ as in [13, Proposition 2.4].
@@ 1. Suppose that $A(\alpha,\beta)$ is a graded down-up algebra
    Other example of a $q^{\prime}$-Heisenberg normal element can be found in [16, Theorem A (d)]. Let $D(v,p)$ be the algebra defined as follows:
    $D(v,p):=k\langle x,y\rangle\left/\middle\langle\begin{aligned} &xy^{2}+vyxy+p^{2}y^{2}x,\\ &x^{3}y+(v+p)x^{2}yx+(p^{2}+pv)xyx^{2}+p^{3}yx^{3}\end{aligned}\right\rangle$
    where $v,p\in k$ and $p\neq 0$. When $v=2p$ holds, an element $x^{2}y+2pxyx+p^{2}yx^{2}\in D(v,p)$ is a $q^{\prime}$-Heisenberg normal element.
@@
    More generally, for a $\mathbb{N}$-graded $k$-algebra AA, let $f$ be a homogeneous element with degree $n$, and $\delta$ a graded derivation of $A$. Suppose that $B=A[X,\delta]$ is a graded Ore extension of $A$ with a degree $1$ element $X$. If $Y=\delta(f)$ is regular in $B[Y]/\langle Y-\delta(f)\rangle$ where $Y$ is a degree $n+1$ element, then $\delta(f)$ is a $q^{\prime}$-Heisenberg normal element of it. Indeed, the definition of the Ore extension gives the following equations;
@@
We will show some basic properties about a $q^{\prime}$-Heisenberg normal element. First, we recall a basic property of a normal element.
Lemma 3.4.
Let $A$ be a $\mathbb{Z}$-graded $k$-algebra, $g$ a regular homogeneous normal element of $A$ with degree $n\geq 1$, and $A^{[n]}$ the $n$-th quasi-Veronese algebra of $A$. Suppose that $\mathbf{g}$ is the element of $A^{[n]}$ defined as
where $\delta_{i,j}$ is the Kronecker delta. Then $\mathbf{g}$ is a regular homogeneous normal element of $A^{[n]}$ with degree $1$.
@@
Let $\nu_{g}$ the graded automorphism determined by $\nu_{g}(a)g=ga$ for $a\in A$. Also, the map defined by $\nu_{\mathbf{g}}^{[n]}((a_{i,j}))=(\nu_{g}(a_{i,j}))$ for $(a_{i,j})\in(A^{[n]})_{r}$ is a graded automorphism of $A^{[n]}$ in [18]. For $\mathbf{a}=(a_{i,j})\in(A^{[n]})_{r}$, the following equation holds;
$\mathbf{ga}=\left(\sum_{l=0}^{n-1}\delta_{l,j}ga_{i,l}\right)=(ga_{i,j})=(\nu_{g}(a_{i,j})g)=\left(\sum_{l=0}^{n-1}\nu_{g}(a_{l,j})\delta_{i,l}g\right)=\nu_{\mathbf{g}}^{[n]}(\mathbf{a})\mathbf{g}.$
Then $\mathbf{g}$ is a regular normal element in $A^{[n]}$ with degree $1$. ∎
Note that, by [3] and [31, Example 2.12], the set $\nu_{\mathbf{g}}^{[n]}=\{(\nu_{\mathbf{g}}^{[n]})^{i}\mid i\in\mathbb{Z}\}$ is a twisting system of $A^{[n]}$ and $\mathbf{g}$ is a central element of the twisted algebra $^{\nu_{\mathbf{g}}^{[n]}}A^{[n]}$.
Remark 3.5.
In the rest of this section, we use the presented notations. For a $\mathbb{Z}$-graded $k$-algebra AA, let $g$ be a regular homogeneous normal element of $A$ with degree $n\geq 1$. We denote an element $\mathbf{g}:=(\delta_{i,j}g)$ of $A^{[n]}$ as $\mathbf{g}$ where $\delta_{i,j}$ is the Kronecker delta. Also, $\nu_{g}$ denotes the graded automorphism of $A$ defined by $\nu_{g}(a)g=ga$ for $a\in A$ and $\nu_{\mathbf{g}}^{[n]}$ denotes the graded automorphism of $A^{[n]}$ defined by $\nu_{\mathbf{g}}^{[n]}((a_{i,j}))=(\nu_{g}(a_{i,j}))$ for $(a_{i,j})\in(A^{[n]})_{r}$.
To investigate a relationship between an algebra having a $q^{\prime}$-Heisenberg normal element and the first Weyl algebra, we recall the definition of the first Weyl algebra $W_{1}(k)$ (see [10]);
$W_{1}(k):=k\langle X,Y\rangle/\langle XY-YX-1\rangle.$
Also, we recall a filtration on the first Weyl algebra called the Bernstein filtration $\mathcal{B}=\{\mathcal{B}_{i}\}$ (see [10]);
$\mathcal{B}_{i}=\begin{cases}\left\{\operatorname*{\mathchoice{\Xdsum}{\Xdsum}{\Xtsum}{\Xtsum}}_{|\alpha|+|\beta|\leq i}c_{\alpha,\beta}Y^{\alpha}X^{\beta}\in W_{1}(k)\middle|c_{\alpha,\beta}\in k\right\}&\text{for}\,i\geq 0;\\ 0&\text{for}\,i<0.\end{cases}$
Proposition 3.6.
[…]
Proof.
We denote the identity of $A^{[n]}$ as $\mathbf{1}$. Let $\varphi:W_{1}(k)\rightarrow\mathcal{R}$ be the $k$-algebra homomorphism defined by mapping of generators as follows;
@@
Since $\nu_{g}(x)=u^{-1}x$, $\nu_{g}(y)=uy$, and $\nu_{g}(g)=g$, the following equations are obtained:
@@
Clearly, the homomorphism $\varphi$ preserves the filtrations $\mathcal{B}$ and $F\mathcal{R}$. ∎
Since the first Weyl algebra $W_{1}(k)$ is simple (see [10, 2.2.1 Theorem]), $\mathcal{R}$ has a subalgebra isomorphic to $W_{1}(k)$ by Proposition 3.6.
Proposition 3.7.
Let $A$ be a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$, g=xyuyxg=xy-uyx a $q^{\prime}$-Heisenberg normal element of $A$ with degree $n$. Then there exists a one-to-one correspondence between the set of isomorphism classes of base-point modules over $A$ and $A/\langle g\rangle$.
Proof.
@@
[…]
$\operatorname{dim}_{k}F_{s}\mathcal{N}=\operatorname{dim}_{k}{}^{\nu_{\mathbf{g}}^{[n]}}(Q(M))_{s}=\operatorname*{\mathchoice{\Xdsum}{\Xdsum}{\Xtsum}{\Xtsum}}_{i=0}^{n-1}\operatorname{dim}_{k}M_{sn-i}=cn\quad\text{ for $s>0$.}$
Hence, there exists a good filtration $\Gamma=\{\Gamma_{i}\}_{i\in\mathbb{Z}}$ of $\mathcal{N}$ and exist $c_{1}\neq 0,c_{0}\in\mathbb{Q}$ such that $\operatorname{dim}_{k}\Gamma_{s}=c_{1}s+c_{0}$ for $s\gg 0$. Furthermore, there exists $s_{0}\in\mathbb{Z}$ such that $\Gamma_{i}\subset F_{i+s_{0}}\mathcal{N}$ for $i\geq 0$ by [10, 8.3.2]. Hence, the inequality $cn=\operatorname{dim}_{k}F_{s+s_{0}}\mathcal{N}\geq\operatorname{dim}_{k}\Gamma_{s}=c_{1}s+c_{0}$ holds for $s\gg 0$. This is a contradiction. Therefore, $M$ is not $g$-torsionfree and then $gM=0$ by [7, Lemma 5]. ∎
@@
Let $M=\bigoplus_{i\in\mathbb{Z}}M_{i}$ be a $g$-tosionfree left bounded graded module such that $\operatorname{dim}_{k}M_{s}=c_{1}s+c_{0}$ for $s\gg 0$ where $c_{1}\neq 0$, $c_{0}\in\mathbb{Q}$ (e.g., $M$ is a line module). By the similar proof of Proposition 3.7, the dehomogenization of $^{\nu_{\mathbf{g}}^{[n]}}Q(M)$ is a holonomic module of $W_{1}(k)$.
4. $q^{\prime}$-Heisenberg normal elements and point modules
In this section, for a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$, we show more details of a relationship between a point scheme of $A$ and a $q^{\prime}$-Heisenberg normal element of $A$. In the end of this section, we show the structure of the set of closed points of the point scheme of the universal enveloping algebra of a color Lie algebra.
4.1. $q^{\prime}$-Heisenberg normal elements and truncated point modules
Throughout this subsection, let $A$ be a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$ unless otherwise specified. In this subsection, we show a relationship between a $q^{\prime}$-Heisenberg normal element $g$ of $A$ and a truncated $g$-torsionfree point module over $A$.
First, we recall useful equations obtained by the definition of a $q^{\prime}$-Heisenberg normal element of $A$. For a $q^{\prime}$-Heisenberg normal element g=xyuyxg=xy-uyx of $A$, the equations $x^{2}y-2uxyx+u^{2}yx^{2}=0$ and $xy^{2}-2uyxy+u^{2}y^{2}x=0$ holds. Then the same equations in a down-up algebra holds as in [12].